# CoxSE: Exploring the Potential of Self-Explaining Neural Networks with Cox Proportional Hazards Model for Survival Analysis

### v1

@@ 4.1 Datasets
Three synthetic datasets are created with different characteristics following the proportional hazards assumption. Moreover, two real datasets were used in the experiments.
4.1.1 Lin Dataset
A synthetic dataset that consists of $50000$ subjects with three features of increasing importance, and a linear relationship between the log-risk $f(\mathbf{x})$ and the predictors as follows:
$f(\mathbf{x})=0.44x_{0}+0.66x_{1}+0.88x_{2}$ (18)
4.1.2 NonLin Dataset
A synthetic dataset that consists of $50000$ subjects with three features and a non-linear relationship between the log-risk $f(\mathbf{x})$ and the predictors and different importance represented by the feature weights as follows:
$f(\mathbf{x})=3x_{0}^{2}+2x_{1}^{2}+x_{2}^{2}$ (19)
4.1.3 NonLinX Dataset
A synthetic dataset that consists of $50000$ subjects with three features and a non-linear relationship between the log-risk $f(\mathbf{x})$ and the predictors, different importance represented by the feature weights, and an interaction between the first two features, as follows:
$f(\mathbf{x})=3x_{0}^{2}+2x_{1}^{2}+x_{2}^{2}+3x_{0}x_{1}$ (20)
4.1.4 FLCHAIN Dataset
This survival analysis benchmark dataset is used to study whether the free light chain (FLC) assay is a good predictor of survival time [29]. The dataset consists of $7874$ subjects with $27$ numerical and one-hot-encoded features, where the FLC is the sum of the two features kappa and lambda. As a result, the study found that a high FLC value was a significant predictor of worse survival.
4.1.5 SEER Dataset
The Surveillance, Epidemiology, and End Results (SEER) database [30], maintained by the US National Cancer Institute, is a comprehensive dataset that gathers detailed information on cancer patients. This dataset includes data on diagnoses, treatments, and survival rates. We utilized version 8.4.3 of the "SEER*Stat" software and the "Incidence – SEER research data, 17 registries" release from November 2022. The dataset encompasses cancer incidences from 2000 to 2020. However, we narrowed our analysis to incidences between 2004 and 2015 to avoid preprocessing challenges due to different coding systems used across various years. The dataset consists of $458117$ subjects with $32$ numerical, label-encoded, and one-hot-encoded features.
5 Experiments
@@ 5.1 Experiment Settings
Five models, all based on Cox Proportional Hazards are compared, CPH, DeepSurv, CoxNAM, CoxSE, and CoxSENAM. All those models are self-explaining except for DeepSurv for which the SHAP method is used to explain. Experiments on the three synthetic and the two real datasets were conducted. Five-fold cross-validation is used to tune the neural networks’ hyperparameters, Number of Layers, Number of Nodes, L2 regularization, Droupout, and Learning rate. In addition to $\alpha$ and $\beta$, the weights of the CoxSE and CoxSENAM loss function. More details about the hyperparameters are available in A.
5.2 Models Performance
Results in Table 1 show that all models performed similarly on the Lin dataset. Also, the four deep-learning-based models performed similarly on the NonLin dataset except for the CPH model, which could not model the non-linear relationship in the data. The more remarkable case is the NonLinX, the non-linear synthetic dataset with interactions between the features. On this dataset, both, DeepSurv and CoxSE performed similarly. However, CoxNAM and CoxSENAM failed to model the feature interaction and converged to a significantly lower C-index. In this case, CPH also failed to model the data and converged to a random model. This difference in performance on the NonLinX dataset is due to the structural design of the NAM-based models CoxNAM and CoxSENAM.
In the cases of the real datasets, all models performed similarly on the FLCHAIN dataset except for the CPH model. More intriguingly, models show three levels of performance on the SEER dataset. The best performance is demonstrated by CoxSE and DeepSurv with no significant difference between the two. On the other hand, CoxNAM and CoxSENAM also performed similarly, however, significantly lower than CoxSE and DeepSurv. Finally, CPH performed significantly lower than the four other models. Such results on the SEER dataset could indicate non-linearity and interaction between the features.
@@ 5.2 Models Performance
CPH 67.67$\pm$0.020 (50.07$\pm$0.078) (50.01$\pm$0.069) (78.33$\pm$0.094) (80.05$\pm$0.103)
DeepSurv 67.68$\pm$0.001 75.30$\pm$0.011 77.92$\pm$0.018 78.45$\pm$0.095 84.35$\pm$0.027
CoxNAM 67.67$\pm$0.011 75.30$\pm$0.012 (70.84$\pm$0.092) 78.59$\pm$0.169 (83.86$\pm$0.018)
CoxSENAM 67.67$\pm$0.012 75.31$\pm$0.006 (70.93$\pm$0.118) 78.51$\pm$0.057 (83.85$\pm$0.029)
CoxSE 67.68$\pm$0.001 75.29$\pm$0.024 77.93$\pm$0.021 78.57$\pm$0.031 84.37$\pm$0.018
Table 1: C-Index performance of all the models on all the datasets. The (.) indicates a significant difference.
@@ 5.3 Aggregated Explanations
The aggregated explanations are computed as the mean of the absolute values of explanations of the explained instances:
$exp_{i}=\frac{1}{n}\sum_{j=0}^{n-1}|f_{exp}^{i}(\mathbf{x}_{j})|$ (21)
where $f_{exp}^{i}(.)$ is the explanation value returned by the model of the feature of index $i$, $\mathbf{x}_{j}$ is the features vector of the instance $j$, and $n$ is the number of the explained instances.
Results illustrated in Figures 45, and 6 show the aggregated explanations of the models on the three synthetic datasets sorted by their average effect. In general, most models agree with the ground truth of the generated data, however, a few intriguing cases deviate from it. Firstly, CoxNAM on the Lin dataset, illustrated in Figure 4(d), ranked the features incorrectly although it achieved the same performance as the other models. This could suggest that modeling the output as an additive function, as in the NAM models, does not guarantee providing correct explanations. Moreover, it indicates the importance of the role of the regularization terms in the CoxSE and CoxSENAM models. It also signifies the proposal of the hybrid model CoxSENAM as a better-regularized version of the NAM-based models. Another case is the CoxNAM explanations on the NonLinX dataset where although it ranked the features correctly, their relative magnitudes are not comparable to the ground truth in contrast to the same-performing model CoxSENAM. Finally, although the CPH model ranked the features correctly on the NonLin dataset and incorrectly on the NonLinX dataset, both explanations can be considered random due to the model’s poor random performance ($\approx 50\%$ C-index) on these two non-linear datasets.
@@ 5.4 Feature Interactions
In the previous section, the results on the NonLinX dataset showed a large discrepancy between Deepsurv and CoxSE on one side and the NAM-based models on the other side. The two NAM-based models, as they have a separate model for each feature, learn the contribution of each feature based only on its corresponding feature neglecting the interactions between the features. This is evident from the experimental results shown in Figure 7. In this experiment, we created seven synthetic datasets based on the log-risk function in Equation 22 with varying the interaction strength factor $\lambda$.
$f(\mathbf{x})=3x_{0}^{2}+x_{1}^{2}+\lambda x_{0}x_{1}$ (22)
DeepSurv model is the most flexible among the other Cox-based models as it holds no design assumptions on the function $f(\mathbf{x})$, other than that it is modeled as a feed-forward neural network. In contrast to DeepSurv, the other models, CPH, CoxSE, CoxNAM, and CoxSENAM force structural constraints making their hypothesis spaces subspaces of the DeepSurvs hypothesis space. Aside from fluctuations resulting from over or under-fitting, DeepSurv performance is guaranteed to be the upper bound for the rest of the Cox-based models. Figure 7 shows the divination of the models’ performance from DeepSurv performance. The NAM-based models’ performance degrades with the increase of the interaction strength, while the CoxSE model maintains an equivalent performance to DeepSurv.
Figure 7: Featue interactions’ effect on models’ performance
5.5 Explanations Stability
The stability is assessed using the Lipschitz constant, Equation 23, which quantifies the model’s sensitivity to a small change in the input [8]. It estimates the maximum change in explanation for a small change in the vicinity of the point of interest.
$L=argmax_{x_{j}\in\epsilon(x_{i})}\frac{\|f_{expl(x_{i})}-f_{expl(x_{j})}\|_{2}}{\|x_{i}-x_{j}\|_{2}}$ (23)
where $\epsilon(x_{i})$ is the neighborhood of $x_{i}$.
@@ 5.5 Explanations Stability
The results illustrated in Figures 8 show that the CPH model has the best stability. This is unsurprising as it is a linear model that has a global explanation for all points. Additionally, CoxSE and CoxSENAM models achieved better stability compared to CoxNAM and DeepSurv(+SHAP). This improvement can be attributed to the regularization term $\mathcal{L}_{2}$ weighted by $\alpha$ in their loss function. In CoxSE and CoxSENAM models, the $\mathcal{L}_{2}$ term controls the model’s flexibility locally. Higher $\alpha$ values result in less flexible models. In the limit, they converge to linear models for very high $\alpha$ values, i.e., the least flexibility and best stability.
@@
Figure 8: Comparison between the models’ stability across different datasets (lower is more stable)
@@ 5.6 Robustness to Non-informative Features
The robustness is computed as the ratio of the informative to the non-informative contributions of the features. The higher the ratio the more robust the model is to noisy features, in which case more contribution comes from informative features relative to non-informative features.
Robustness $=$ $\sum_{i=0}^{N}\frac{y_{i}^{\Omega}}{y_{i}^{\overline{\Omega}}}$ (24)
@@
$y^{\Omega}$ $=$ $\sum_{x_{i}\in\Omega(\mathbf{x})}w_{i}(\mathbf{x})x_{i}$ (25)
$y^{\overline{\Omega}}$ $=$ $\sum_{x_{i}\in\overline{\Omega(\mathbf{x})}}w_{i}(\mathbf{x})x_{i}$ (26)
where $\Omega(\mathbf{x})$ is the set of informative features of $\mathbf{x}$ and $\overline{\Omega(\mathbf{x})}$ is its complement, i.e., the set of non-informative features.
In this experiment, to guarantee that the three models, CoxSE, CoxNAM, and CoxSENAM can model the data with comparable performances, we created synthetic datasets with non-linear log-risk as in Equation 22 without interaction ($\lambda=0$ ) varying the dataset size from $1$k to $50$k samples. Each dataset contains $10$ features, two of which are informative and the rest are non-informative features. To study the models’ robustness and the effect of dataset size and the regularization parameters, we ran several experiments varying the dataset size and $\alpha$ and $\beta$ for the CoxSE and the CoxSENAM models.
@@
The results in Figure 9 show that all models benefit from increasing the dataset size in improving the C-index, Figure 9(b) and models’ robustness, Figure 9(a). Moreover, for very small and very large dataset sizes, CoxSE and CoxNAM have a comparable robustness score. At the same time, CoxNAM Shows better robustness than CoxSE in med-size datasets as illustrated in Figure 9(a). This can be attributed to the design structure of CoxNAM where non-informative features only contribute to the final decision in the neural network in the form of additive noise. In CoxSE, however, computing feature relevances incorporates all the informative and non-informative features passed through the neural network in a complicated way. This makes it harder for the CoxSE model to isolate the noise. However, the hybrid model, CoxSENAM demonstrated better robustness which can be attributed to the extra regularization term $\mathcal{L}_{3}$ weighted by $\beta$ in its loss function. To illustrate that, Figure 10, shows the effect of dataset size and the change in $\beta$, the weight of the $\mathcal{L}_{3}$ term common for both CoxSE and CoxSENAM loss functions. Increasing $\beta$ does not seem to have any effect on the CoxSE model’s robustness, however, the effect is significant in the case of CoxSENAM as shown in Figures 10(a) and 10(b) respectively.
@@
(b) CoxSENAM: Robustness
(c) CoxNAM: Robustness
(d) CoxSE: C-Index
(e) CoxSENAM: C-Index
(f) CoxNAM: -Index

### v2

@@ 5.1 Datasets
Three synthetic datasets are created with different characteristics based on the proportional hazards assumption following the simulations in [Kvamme:2019]. Moreover, two real datasets were used in the experiments.
Lin Dataset: A synthetic dataset that consists of $50,000$ subjects with three features of increasing importance, and a linear relationship between the log-risk $f(\bm{x})$ and the predictors as follows:
$f(\bm{x})=0.44x_{0}+0.66x_{1}+0.88x_{2}$ (19)
NonLin Dataset: A synthetic dataset that consists of $50,000$ subjects with three features and a non-linear relationship between the log-risk $f(\bm{x})$ and the predictors, with different importance represented by the feature weights as follows:
$f(\bm{x})=3x_{0}^{2}+2x_{1}^{2}+x_{2}^{2}$ (20)
NonLinX Dataset: A synthetic dataset that consists of $50,000$ subjects with three features and a non-linear relationship between the log-risk $f(\bm{x})$ and the predictors, different importance represented by the feature weights, and an interaction between the first two features, as follows:
$f(\bm{x})=3x_{0}^{2}+2x_{1}^{2}+x_{2}^{2}+3x_{0}x_{1}$ (21)
FLCHAIN Dataset: This survival analysis benchmark dataset is used to study whether the free light chain (FLC) assay is a good predictor of survival time [Dispenzieri:2012]. The dataset consists of $7,874$ subjects with $27$ numerical and one-hot-encoded features, where the FLC is the sum of the two features kappa and lambda. As a result, the study found that a high FLC value was a significant predictor of worse survival.
SEER Dataset: The Surveillance, Epidemiology, and End Results (SEER) database [seer:2016], maintained by the US National Cancer Institute, is a comprehensive dataset that gathers detailed information on cancer patients. This dataset includes data on diagnoses, treatments, and survival rates. We utilized version 8.4.3 of the "SEER*Stat" software and the "Incidence – SEER research data, 17 registries" release from November 2022. The dataset encompasses cancer incidences from 2000 to 2020. However, we narrowed our analysis to incidences between 2004 and 2015 to avoid preprocessing challenges due to different coding systems used across various years. The dataset consists of $458,117$ subjects with $32$ numerical, label-encoded, and one-hot-encoded features.
@@ 5.2 Experiment Settings
Five models, all based on Cox Proportional Hazards, are compared: CPH, DeepSurv, CoxNAM, CoxSE, and CoxSENAM. All those models are self-explaining except for DeepSurv, for which the SHAP method is used to explain. Experiments on the three synthetic and the two real datasets were conducted. To provide benchmarking context, we also report results from two non-Cox approaches, Random Survival Forest (RSF) [Ishwaran_2008] and DeepHit [Lee_2018]. However, as our focus is on interpretability within Cox-based frameworks, these baselines are not included in the detailed discussion. Five-fold cross-validation is used to tune the models. More details about the hyperparameters are provided in A.
5.3 Performance of the Models
The results in Table 1 show that all models performed similarly on the Lin dataset. Also, the four deep-learning-based models performed similarly on the NonLin dataset except for the CPH model, which could not model the non-linear relationship in the data. The more remarkable case is the NonLinX, the non-linear synthetic dataset with interactions between the features. On this dataset, both DeepSurv and CoxSE performed similarly. However, CoxNAM and CoxSENAM failed to model the feature interaction and converged to a significantly lower C-index. In this case, CPH also failed to model the data and converged to a random model. This difference in performance on the NonLinX dataset is due to the structural design of the NAM-based models, CoxNAM and CoxSENAM, which cannot handle feature interactions.
In the cases of the real datasets, all models except CPH performed similarly on the FLCHAIN dataset. Intriguingly, the models show three levels of performance on the SEER dataset. The best performance is demonstrated by CoxSE and DeepSurv, with no significant difference between them. On the other hand, CoxNAM and CoxSENAM also performed similarly, but significantly below CoxSE and DeepSurv. Finally, CPH performed significantly lower than the four other models. Such results on the SEER dataset could indicate non-linearity and interaction between the features.
Table 1: Concordance Index (C-Index) performance of all models across all datasets. Values in parentheses (.) denote statistically significant differences based on a paired t-test over 5-fold cross-validation.
@@ 5.3 Performance of the Models
[…]
Reference baselines (non-Cox) models
[…]
Building upon the previous analysis of model performance, we further report the average execution time per run for each model in Table 2. The results indicate that CoxSE, CoxSENAM, CoxNAM, and DeepSurv exhibit similar execution times across the majority of datasets. Notably, more substantial differences were observed on the SEER dataset—the largest in the study—where CoxSENAM required the longest time to complete model fitting.
Table 2: Average execution time (in seconds) of the models’ 5-fold runs on each dataset.
[…]
Reference baselines (non-Cox) models
[…]
@@ 5.4 Aggregated Explanations
The aggregated explanations are computed as the mean of the absolute values of the explanations of the explained subjects:
$F_{expl}^{i}=\frac{1}{n}\sum_{j=0}^{n-1}|f_{expl}^{i}(\bm{x}_{j})|$ (22)
where $f_{expl}^{i}(.)$ is the local explanation value returned by the model of the feature of index $i$, $\bm{x}_{j}$ is the feature vector of the subject $j$, and $n$ is the number of explained subjects. In particular, these local explanations are $w_{i}$ for the CPH model (Eq. 6), $w_{\theta_{i}}(\bm{x})$ for CoxSE (Eq. 13), $w_{\theta_{i}}(x_{i})$ for CoxSENAM (Eq. 14), $g_{i}(x_{i})$ for CoxNAM (Eq. 9), and Shapley values for DeepSurv.
The results illustrated in Figures 45, and 6 show the aggregated explanations of the models on the three synthetic datasets sorted by their average effect. In general, most models agree with the ground truth of the generated data, however, a few intriguing cases deviate from it. Firstly, CoxNAM on the Lin dataset, illustrated in Figure 4(d), ranked the features incorrectly, although it achieved the same performance as the other models. This could suggest that modeling the output as an additive function, as in the NAM models, does not guarantee providing correct explanations. Moreover, it indicates the importance of the role of the regularization terms in the CoxSE and CoxSENAM models. It also signifies the proposal of the hybrid model CoxSENAM as a better-regularized version of the NAM-based models. Another case is the CoxNAM explanations on the NonLinX dataset (Figure 6(d)), where, although it ranked the features correctly, their relative magnitudes are not comparable to the ground truth in contrast to the same-performing model CoxSENAM (Figure 6(b)). Finally, although the CPH model ranked the features correctly on the NonLin dataset and incorrectly on the NonLinX dataset, both explanations can be considered random due to the model’s poor random performance ($\approx 50\%$ C-index) on these two non-linear datasets.
@@ 5.5 Feature Interactions
In the previous section, the results on the NonLinX dataset showed a large discrepancy between Deepsurv and CoxSE on one side and the NAM-based models on the other side. The two NAM-based models, as they have a separate model for each feature, learn the contribution of each feature based only on its corresponding feature, neglecting the interactions between the features. This is evident from the experimental results shown in Figure 7. In this experiment, we created seven synthetic datasets based on the log-risk function in Eq. 23 by varying the interaction strength factor $\lambda$.
$f(\bm{x})=3x_{0}^{2}+2x_{1}^{2}+x_{2}^{2}+\lambda x_{0}x_{1}$ (23)
DeepSurv model is the most flexible among the Cox-based models considered, as it imposes no structural assumptions on the function $f(\bm{x})$, beyond being represented by a feed-forward neural network. In contrast, CPH, CoxSE, CoxNAM, and CoxSENAM force structural constraints, making their hypothesis spaces subspaces of DeepSurv’s. While this flexibility allows DeepSurv to approximate more complex relationships, its performance is still influenced by several factors, including the amount and quality of available data, the effectiveness of regularization, and optimization challenges during training. As a result, superior performance is not guaranteed in every setting. Figure 7 shows the deviation of the models’ performance from DeepSurv performance. The NAM-based models’ performance degrades with the increase of the interaction strength, while the CoxSE model maintains an equivalent performance to DeepSurv.
Figure 7: Impact of feature interactions on model performance - as a deviation from the DeepSurv performance
5.6 Stability of Explanations
The stability is assessed using the Lipschitz constant, Eq. 24, which quantifies the model’s sensitivity to a small change in the input [Alvarez2018]. It estimates the maximum change in explanation for a small change in the vicinity of the point of interest.
$L=argmax_{\bm{x}_{j}\in\epsilon(\bm{x}_{i})}\frac{\|f_{expl(\bm{x}_{i})}-f_{expl(\bm{x}_{j})}\|_{2}}{\|\bm{x}_{i}-\bm{x}_{j}\|_{2}}$ (24)
where $\epsilon(\bm{x}_{i})$ is the neighborhood of $\bm{x}_{i}$.
@@ 5.6 Stability of Explanations
The results illustrated in Figure 8 show that the CPH model has the best stability. This is unsurprising as it is a linear model that has a global explanation for all points. Additionally, CoxSE and CoxSENAM models achieved better stability compared to CoxNAM and DeepSurv(+SHAP). This improvement can be attributed to the regularization term $\mathcal{L}_{2}$ weighted by $\alpha$ in their loss functions. In CoxSE and CoxSENAM models, the $\mathcal{L}_{2}$ term controls the model’s flexibility locally. Higher $\alpha$ values result in less flexible models. In the limit, they converge to linear models for very high $\alpha$ values, i.e., the least flexibility and best stability.
@@
Figure 8: Comparison of model stability across datasets, where lower Lipschitz estimates indicate greater stability
To dig deeper into the effect of the two hyperparameters $\alpha$ and $\beta$ on the stability and performance, we created synthetic datasets with non-linear log-risk as in Eq. 23 without interaction ($\lambda=0$ ). Consequently, we trained CoxSE and CoxSENAM models varying $\alpha$ and $\beta$. The results in Figure 9 show that increasing $\alpha$ increases the stability as shown by the decline of the Lipschitz estimate in Figure 9(a). However, very high values of $\alpha$ limit the flexibility of the model, which causes a decline in the performance as shown in Figure 9(b). More importantly, for the CoxSE model, increasing $\alpha$ to a certain degree increased the agreement between the ground truth explanations and the predicted ones, as shown in Figure 9(c). However, after a certain degree, as the performance drops, the agreement with the ground truth explanations also drops due to underfitting.
On the other hand, increasing $\beta$ negatively affects the stability, Figure 9(d), which can be attributed to the rigidity induced by the $\mathcal{L}_{3}$ term in the loss function that favors more sparse learned weights. Moreover, very high values of $\beta$ push all the learned weights towards zero, enhancing the stability; however, this causes a collapse in the model’s performance and the correlation with the ground truth explanations, as illustrated in Figures 9(d)9(e), and 9(f).
(a) Stability vs. alpha
(b) C-index vs. alpha
(c) Correlation vs. alpha
(d) Stability vs. beta
(e) C-index vs. beta
(f) Correlation vs. beta
Figure 9: The effect of $\alpha$ and $\beta$ on the stability, performance, and rank correlation with ground truth explanations
As demonstrated in the previous analysis, varying $\alpha$ and $\beta$ impacts both the model’s stability, predictive performance, and alignment with ground truth explanations. Building on these findings, the analysis suggests that optimal values for $\alpha$ and $\beta$ can be selected by tuning the model to achieve peak performance, while choosing the highest possible regularization strengths that do not compromise that performance.
@@ 5.7 Robustness to Non-informative Features
The robustness is computed as the ratio of the informative to the non-informative contributions of the features. The higher the ratio, the more robust the model is to noisy features, in which case more contribution comes from informative features relative to non-informative features.
Robustness $=$ $\frac{1}{n}\sum_{i=0}^{n-1}\frac{y_{i}(\Omega)}{y_{i}(\overline{\Omega})}$ (25)
@@
$y(\Psi)$ $=$ $\sum_{j\in\Psi}|C_{j}(\bm{x})|$ (26)
where $\Psi$ represents a subset of the features, $\Omega$ is the set of informative features of $\bm{x}$, $\overline{\Omega}$ is its complement, i.e., the set of non-informative features, and $n$ the number of studied subjects. $C_{j}(\bm{x})$ is the contribution of the feature $x_{j}$ in the in the total output $y$, where it is ($C_{j}(\bm{x})=w_{j}(\bm{x})x_{j}$) for CoxSE, ($C_{j}(\bm{x})=w_{j}(x_{j})x_{j}$) for CoxSENAM, and ($C_{j}(\bm{x})=f_{j}(x_{j})$) for CoxNAM.
In this experiment, to guarantee that the three models, CoxSE, CoxNAM, and CoxSENAM, can model the data with comparable performances, we created synthetic datasets with non-linear log-risk as in Eq. 23 without interaction ($\lambda=0$ ), varying the dataset size from 11k to 5050k samples. Each dataset contains $10$ features, two of which are informative and the rest are non-informative features. To study the models’ robustness and the effect of dataset size and the regularization parameters, we ran several experiments varying the dataset size. We also varied $\alpha$ and $\beta$ for the CoxSE and the CoxSENAM models.
@@
The results in Figure 10 show that all models benefit from increasing the dataset size in improving the C-index (Figure 10(b)) and models’ robustness (Figure 10(a)). Moreover, for very small and very large dataset sizes, CoxSE and CoxNAM have a comparable robustness score. At the same time, CoxNAM shows better robustness than CoxSE in med-size datasets as illustrated in Figure 10(a). This can be attributed to the design structure of CoxNAM, where non-informative features only contribute to the final decision in the neural network in the form of additive noise. In CoxSE, however, computing feature relevances incorporates all the informative and non-informative features passed through the neural network in a complicated way. This makes it harder for the CoxSE model to isolate the noise. However, the hybrid model, CoxSENAM demonstrated better robustness, which can be attributed to the extra regularization term $\mathcal{L}_{3}$ weighted by $\beta$ in its loss function. To illustrate this, Figure 11 shows the effect of dataset size and the change in $\beta$, the weight of the $\mathcal{L}_{3}$ term common for both CoxSE and CoxSENAM loss functions. Increasing $\beta$ does not seem to have any effect on the CoxSE model’s robustness, however, the effect is significant in the case of CoxSENAM, as shown in Figures 11(a) and 11(c) respectively.
@@
(b) CoxSE: C-index
(c) CoxSENAM: Robustness
(d) CoxSENAM: C-index
@@
On the other hand, increasing $\alpha$ doesn’t affect the robustness of CoxSE and CoxSENAM models. However, for very high values of $\alpha$, the models’ performance drops significantly due to underfitting, which in turn leads to a decline in the CoxSENAM robustness, as illustrated in Figure 12.
(a) CoxSE: Robustness
(b) CoxSE: C-index
(c) CoxSENAM: Robustness
(d) CoxSENAM: C-index
Figure 12: Comparison between the models’ performance and robustness across different dataset sizes and different $\alpha$ values